Question ? How can you tell by looking at a quadratic equation whether it will have a maximum or a minimum?
Thank You

Answer by nabla(475)

If the coefficient in front of the x^2 term is positive, the quadratic equation will have a minimum. If the coefficient is negative, the quadratic equation will have a maximum. This is due to the square term having the most influence in the resultant values for the function.
We can apply calculus to ax^2+bx+c=f(x)
f'(x)=2ax+b
So, 2ax+b=0 will be the min or max.
This happens at x=-b/2a . This is the formula you will find in your text for the location of the vertex of a quadratic equation.